A tennis ball is thrown upwards with sufficient speed to reach a height of 12 m.


(i) What is the kinetic energy of the tennis ball at a height of 12 m?


(ii) Calculate the speed of the tennis ball when it reaches a height of 6 m.


1
Expert's answer
2022-04-15T16:40:38-0400

(i) What is the kinetic energy of the tennis ball at a height of 12 m?

Answer: zero

(ii) Calculate the speed of the tennis ball when it reaches a height of 6 m.

"\frac{mv^2}{2}+mgh_2=mgh_1"

"v=\sqrt{2g(h_1-h_2)}"

"v=\sqrt{2*9.8(12-6)}=11\:\rm m/s"